OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–4, –3) and (12, 1) is y – 1 = (x – 12). What is the standard form of the equation for this line? (A) x – 4y = 8 (B) x – 4y = 2 (C) 4x – y = 8 (D) 4x – y = 2

OpenStudy (anonymous):

?

OpenStudy (welshfella):

slope of the line = ( 1 - (-3) / 12 - (-4) = 4/16 = 1/4 plugging in y = -3 , x = -4 and m = 1/4 into the point-slope form:- y - (-3) = (1/4)(x - (-4) y + 3 = (1/4) x + 1 multiply through by 4 4y + 12 = x + 4 can you continue/

OpenStudy (anonymous):

ima try to

OpenStudy (welshfella):

subtract 4 from both sides 4y + 8 = x subtract 4y from both sides 8 = x - 4y which is the same as x - 4y = 8 Standard form

OpenStudy (welshfella):

the last step was just swapping sides

OpenStudy (anonymous):

okay i got it thank you?